OpenStudy (anonymous):

multiply. rewrite the fraction in lowest terms (3x/4)(7/xy)

OpenStudy (lalaly):

\[\frac{3\cancel{x}}{4} \times \frac{7}{\cancel{x}y} = \frac{21}{4y}\]